To prepare two liters of a 10% solution from a pure drug, you would need ____?
Answer by stanbon(75887)

To prepare two liters of a 10% solution from a pure drug, you would need ____?
---------------------------------
Equation:
Let x be the amount of water you add.
pure drug + water = 10% mix
1.00(2-x) + 0(x) = 0.10*2
Multiply thru by 100 to get:
100*2 - 100x + 0 = 10*2
-100x = -90*2
x = 1.8 liters (amt. of water needed for the mixture
----
2-x = 0.2 liters (amt. of pure drug needed for the mixture
==================
Cheers,
Stan H.